# Multi-objective Bayesian Optimisation of Spinodoid Cellular Structures for Crush Energy Absorption

### v1

@@ 1 Introduction
Building on the concept of data-driven optimisation, structures can be inverse-designed by starting with desired performance characteristics and working backwards to generate optimal designs. This approach leverages machine learning techniques such as neural networks to capture complex, non-linear relationships [50], convolutional neural networks (CNNs) for handling spatially varying designs, and generative adversarial networks (GANs) for exploring complex design spaces [51, 52]. However, many of these methods are highly data-dependent, requiring large, high-quality datasets that can be costly to generate, making it difficult to acquire the necessary complex nonlinear data due to computational demands. To mitigate this challenge, physics-informed machine learning has been implemented, where experimental data is incorporated to tailor complex structures for large-deformation scenarios [53].
This challenge becomes even more pronounced in real-world applications, such as automotive crashworthiness or aerospace structures, where focusing solely on single-objective optimization—such as independently maximizing energy absorption (EA) or minimizing peak force (PF)—is often inadequate. A balance between multiple conflicting objectives must be achieved simultaneously. In response to these multifaceted optimisation problems, researchers have turned to multi-objective optimisation (MOO) techniques, allowing the identification of optimal solutions that meet diverse requirements. Past implementations to solve such a problem included the use of evolutionary algorithms [54] such as the Non-dominated Sorting Genetic Algorithm (NSGA-II), coupled with an external surrogate model, such as the Radial Basis Function [55], polynomial response surface method [56], and the use of Gaussian Processes (or kriging) [57, 58]. Other implementations include the use of particle swarm-based optimisation [59], and hybrid approaches using the VIKOR method [60, 61]. Recently, neural networks have also been utilised as surrogates for MOO problems [62, 63].
@@
Building on these advantages, multi-objective Bayesian Optimisation (MOBO) extends the capabilities of BO to handle multiple conflicting objectives. The current advancements in MOBO focus on improving scalability and efficiency in high-dimensional spaces by enhancing sampling strategies [75]. These include leveraging deep Gaussian Processes (DGPs), which stack multiple Gaussian Processes to better capture complex dependencies in the data [76], and combining Bayesian statistics with neural networks to form Bayesian Neural Networks (BNNs), which provide scalable and flexible surrogate models [77]. Furthermore, enhanced acquisition functions, such as Expected Hypervolume Improvement (EHI) and Predictive Entropy Search for Multi-Objective Optimization (PESMO), along with their advanced extensions [78, 79, 80], have significantly improved the ability to balance exploration of the search space and exploitation of known promising regions, thereby accelerating convergence to the Pareto front and enabling efficient design space exploration.
@@
However, many real-world problems involve non-linear, non-convex constraints, including manufacturing limitations or safety criteria. Existing MOBO frameworks often struggle to effectively incorporate or optimise under these constraints. To address this, our research introduces a gradient-based filtering approach within the MOBO process to identify and remove structures prone to densification. By dynamically filtering out infeasible solutions early, the framework ensures convergence to robust, manufacturable designs while reducing computational cost. Building on this, the proposed framework leverages MOBO to design spinodoid topologies that incorporate elastic-plastic and fracture material behaviours. Data is generated using finite element method (FEM) simulations, which predict post-yield performance metrics such as PF, and EA. This enables a comprehensive evaluation of the structural response and facilitates the optimisation of design parameters to enhance EA while reducing PF.
@@
The novelty of this study lies in three key advancements, offering a comprehensive solution for optimising energy-absorbing structures: (1) The proposed optimisation framework significantly enhances MOBO by efficiently handling multiple conflicting objectives, enabling comprehensive exploration of the design space and balancing trade-offs that traditional methods often struggle to address. Leveraging pre-existing data reduces computational costs, requiring fewer simulations and achieving faster convergence. Additionally, the framework incorporates real-world constraints, such as manufacturing tolerances and fabrication defects by integrating complex material constitutive behaviours into FEM simulations, leading to robust and reliable designs. The introduction of a gradient-based filtering mechanism to identify and exclude parameter sets that result in structures that undergo densification which may degrade the performance of designs. (2) This study applies MOBO to optimise structures exhibiting highly non-linear crush behaviours, a challenging domain that has not been extensively explored. By using FEM simulations to evaluate elastic-plastic and fracture material responses, the framework optimises critical performance metrics, such as PF and EA. (3) The study pioneers the application of spinodoid topologies for crush energy absorption, demonstrating their tunability of anisotropy and energy absorption. This novel use of spinodoids leverages their unique structural characteristics to enhance energy-absorbing capabilities, making them particularly suitable for crashworthiness applications. The strategy opens up new possibilities for utilising spinodoid designs in diverse engineering disciplines.
@@ 2.1 Design parameters of Spinodoid cellular structures
To illustrate the influence of the three cone angles $\theta_{i}^{i=1,2,3}$, four distinct structures were created, each showcasing a unique combination of these angles. This approach demonstrates the extensive parametric design space offered by spinodoid topologies, with these structures being depicted in Fig. 2. The design space is capable of producing anisotropic structures, as exemplified by Fig. 2(a) and (d). These structures are generated as a result of imposing small cone angle values, with at least one angle being set to zero. These configurations exhibit stiffness favouring at least one of the principal directions. For instance, the ‘Columnar’ structure in Fig. 2(a) arises from a choice of two small, non-zero angle values, which results in a network of interconnected pillar-like structures. Another form of anisotropic structure, termed ‘Lamellar’, is demonstrated by Fig. 2(d), where one non-zero cone angle value produces a structure featuring several interconnected layers, with the layer count being proportional to the wave number. Demonstration of controlling the anisotropy of structures through a choice of cone angles can be found in A.1. Conversely, Fig. 2(c) showcases the formation of a ‘Cubic’ topology through the use of three non-zero cone angles of equivalent magnitude, which produces a structure with identical stiffnesses in all principal directions. Meanwhile, ‘Isotropic’ structures, as shown in Fig. 2(b), emerge from any combination of angles featuring sufficiently large values of cone angles.
@@ 2.2 FEM setup and material characterisation
The spinodoid topologies of dimensions $L\times L\times L$ (where $L=40$ mm) were generated using MATLAB, and ABAQUS Explicit solver was utilised for FEM analysis, with tetrahedral element types used for meshing the structures. The performance of different topologies, generated with varying design parameters, was evaluated through these FEM simulations. The simulation setup involved using two rigid rectangular plates, one acting as the loader, moving until a maximum strain of 50% was reached, and the other being the anvil, both placed at the top and bottom of the structure, respectively. Furthermore, general contact was defined to simulate the contact between the platens as well as generating self-contact to avoid self-intersection, allowing densification to occur. For validation, spinodoid structures were fabricated using additive manufacturing, inherently introducing directional dependency within the structure. To capture this within the simulations, engineering constants were utilised to describe the linear elasticity in orthotopic materials, with three moduli in each principal direction $E_{1},E_{2},E_{3}$; Poisson’s ratio $\nu_{12},\nu_{13},\nu_{23}$; and the shear moduli $G_{12},G_{13},G_{23}$. These constants are then used to construct the elastic compliance, which is as follows
@@
$F(\sigma_{22}-\sigma_{33})^{2}+G(\sigma_{33}-\sigma_{11})^{2}+H(\sigma_{11}-\sigma_{22})^{2}+2L\sigma^{2}_{23}+2M\sigma^{2}_{31}+2N\sigma^{2}_{23}=1$ (5)
$F=\frac{1}{2}\left[\frac{1}{(\sigma_{2}^{y})^{2}}+\frac{1}{(\sigma_{3}^{y})^{2}}-\frac{1}{(\sigma_{1}^{y})^{2}}\right]\qquad L=\frac{1}{2(\tau^{y}_{23})^{2}}$ (6)
$G=\frac{1}{2}\left[\frac{1}{(\sigma_{3}^{y})^{2}}+\frac{1}{(\sigma_{1}^{y})^{2}}-\frac{1}{(\sigma_{2}^{y})^{2}}\right]\qquad M=\frac{1}{2(\tau^{y}_{31})^{2}}$
$H=\frac{1}{2}\left[\frac{1}{(\sigma_{1}^{y})^{2}}+\frac{1}{(\sigma_{2}^{y})^{2}}-\frac{1}{(\sigma_{3}^{y})^{2}}\right]\qquad N=\frac{1}{2(\tau^{y}_{12})^{2}}$
where $\sigma_{1}^{y}$, $\sigma_{2}^{y}$ and $\sigma^{y}_{3}$ are the normal yield stresses with respect to the axes of anisotropy and $\tau_{12}^{y}$, $\tau_{23}^{y}$ and $\tau_{31}^{y}$ are the yield stresses in shear, obtained from experiments.
@@
Tensile and compressive tests were conducted on coupons, following the ISO-527-1’s test specimen dimensions, and on cubes of dimensions $20\times 20\times 20$ $\text{mm}^{3}$. It should be noted that here, a relatively small cube size was used to ensure a high-capacity load cell was not required to crush the sample. The test specimens were fabricated using fused deposition modelling with the UltiMaker S5, with samples being printed with three different printing directions: $0^{\circ}$, $45^{\circ}$, and $90^{\circ}$. The results from the material characterisation tests have been plotted in Fig. 3. Both quasi-static compression and tensile tests were conducted at a strain rate of $0.1~{}s^{-1}$, using an INSTRON-5967 machine. The results of these tests have been graphed in Fig. 3. Material properties were subsequently extracted and used to apply the material constitutive model to the spinodoid structures.
@@ 2.3 Experimental validation
To validate the FEM model, the four spinodoid structures described in Fig. 2 were employed. Similar to the testing samples, the spinodoid structures of dimensions $20\times 20\times 20$ mm3 were additively manufactured using UltiMaker S5. Furthermore, the tests were conducted on samples with a relative density of 0.3 and a wave number of $24\pi$. A smaller dimension of spinodoid structures was chosen, along with low relative density to offset the densification, and by extension, not require a higher capacity load cell. Compressive testing was conducted using a quasi-static strain rate of 0.1 $s^{-1}$. This strain rate was carefully chosen to closely match the conditions of the material characterisation test, minimising the influence of strain-rate variations on the results. To ensure the separation of scales, a larger wave number was utilised as a result of using a smaller-sized structure, wave numbers larger than $24\pi$ would result in geometries with features that cannot be resolved by the 3D printer, even with layer height set to a minimal. Fig. 4 shows the results of the compressive testing of the spinodoid structures. The compressive tests were conducted on three samples for each type of topology to ensure repeatability as well as to consider the variation in crushing behaviour resulting from manufacturing defects. In the figure, the black lines show the experimental results, while the red lines show the results extracted from the FEM simulations. Directly using the material constants obtained from the material characterisation tests caused an overprescription of stiffness within the structures, resulting in a mismatch between experimental and simulation results. This was remedied by reducing the material constants to values that were 20% of the original. Introducing this artificial reduction in the stiffness constants accounts for any structural defects and artefacts created during the manufacturing of the prototypes, which may weaken the structural integrity of the structures. However, for ‘Lamellar’ structure, increased stiffness values were utilised with a magnitude of 60% of the originally extracted stiffness from the material characterisation tests. Here, ‘Lamellar’-type structures are topologies with layers that are perpendicular to the loading direction. This reduced stiffness penalty was applied to structures with $0\leq\theta_{1,2}\leq 5$ and $0\leq\theta_{3}\leq 45$.

### v2

@@ 1 Introduction
Building on the concept of data-driven optimisation, structures can be inverse-designed by starting with desired performance characteristics and working backwards to generate optimal designs. This approach leverages machine learning techniques such as neural networks to capture complex, non-linear relationships [50], convolutional neural networks (CNNs) for handling spatially varying designs, and generative adversarial networks (GANs) for exploring complex design spaces [51, 52]. However, many of these methods are highly data-dependent, requiring large, high-quality datasets that can be costly to generate, making it difficult to acquire the necessary complex non-linear data due to computational demands. To mitigate this challenge, physics-informed machine learning has been implemented, where experimental data is incorporated to tailor complex structures for large-deformation scenarios [53].
This challenge becomes even more pronounced in real-world applications, such as automotive crashworthiness or aerospace structures, where focusing solely on single-objective optimisation—such as independently maximising energy absorption (EA) or minimising peak force (PF)—is often inadequate. A balance between multiple conflicting objectives must be achieved simultaneously. In response to these multifaceted optimisation problems, researchers have turned to multi-objective optimisation (MOO) techniques, allowing the identification of optimal solutions that meet diverse requirements. Past implementations to solve such a problem included the use of evolutionary algorithms [54] such as the Non-dominated Sorting Genetic Algorithm (NSGA-II), coupled with an external surrogate model, such as the Radial Basis Function [55], polynomial response surface method [56], and the use of Gaussian Processes (or kriging) [57, 58]. Other implementations include the use of particle swarm-based optimisation [59], and hybrid approaches using the VIKOR method [60, 61]. Recently, neural networks have also been utilised as surrogates for MOO problems [62, 63].
@@
Building on these advantages, multi-objective Bayesian Optimisation (MOBO) extends the capabilities of BO to handle multiple conflicting objectives. The current advancements in MOBO focus on improving scalability and efficiency in high-dimensional spaces by enhancing sampling strategies [75]. These include leveraging deep Gaussian Processes (DGPs), which stack multiple Gaussian Processes to better capture complex dependencies in the data [76], and combining Bayesian statistics with neural networks to form Bayesian Neural Networks (BNNs), which provide scalable and flexible surrogate models [77]. Furthermore, enhanced acquisition functions, such as Expected Hypervolume Improvement (EHI) and Predictive Entropy Search for Multi-Objective Optimisation (PESMO), along with their advanced extensions [78, 79, 80], have significantly improved the ability to balance exploration of the search space and exploitation of known promising regions, thereby accelerating convergence to the Pareto front and enabling efficient design space exploration.
@@
However, many real-world problems involve non-linear, non-convex constraints, including manufacturing limitations or safety criteria. Existing MOBO frameworks often struggle to effectively incorporate or optimise under these constraints. To address this, our research introduces a gradient-based filtering approach within the MOBO process to identify and remove structures prone to densification. By dynamically filtering out infeasible solutions early, the framework ensures convergence to robust, manufacturable designs while reducing computational cost. Building on this, the proposed framework leverages MOBO to design spinodoid topologies that incorporate elastic-plastic material behaviours. Data is generated using finite element method (FEM) simulations, which predict post-yield performance metrics such as PF, and EA. This enables a comprehensive evaluation of the structural response and facilitates the optimisation of design parameters to enhance EA while reducing PF.
@@
The novelty of this study lies in three key advancements, offering a comprehensive solution for optimising energy-absorbing structures: (1) The proposed optimisation framework significantly enhances MOBO by efficiently handling multiple conflicting objectives, enabling comprehensive exploration of the design space and balancing trade-offs that traditional methods often struggle to address. Leveraging pre-existing data reduces computational costs, requiring fewer simulations and achieving faster convergence. Additionally, the framework incorporates real-world constraints, such as manufacturing tolerances and fabrication defects by integrating complex material constitutive behaviours into FEM simulations, leading to robust and reliable designs. The introduction of a gradient-based filtering mechanism to identify and exclude parameter sets that result in structures that undergo densification which may degrade the performance of designs. (2) This study applies MOBO to optimise structures exhibiting highly non-linear crush behaviours, a challenging domain that has not been extensively explored. By using FEM simulations to evaluate elastic-plastic material responses, the framework optimises critical performance metrics, such as PF and EA. (3) The study pioneers the application of spinodoid topologies for crush energy absorption, demonstrating their tunability of anisotropy and energy absorption. This novel use of spinodoids leverages their unique structural characteristics to enhance energy-absorbing capabilities, making them particularly suitable for crashworthiness applications. The strategy opens up new possibilities for utilising spinodoid designs in diverse engineering disciplines.
@@ 2.1 Design parameters of Spinodoid cellular structures
To illustrate the influence of the three cone angles $\theta_{i}^{i=1,2,3}$, four distinct structures were created, each showcasing a unique combination of these angles. This approach demonstrates the extensive parametric design space offered by spinodoid topologies, with these structures being depicted in Fig. 2. The design space is capable of producing anisotropic structures, as exemplified by Fig. 2(a) and (d). These structures are generated as a result of imposing small cone angle values, with at least one angle being set to zero. These configurations exhibit stiffness favouring at least one of the principal directions. For instance, the ‘Columnar’ structure in Fig. 2(a) arises from a choice of two small, non-zero angle values, which results in a network of interconnected pillar-like structures. Another form of anisotropic structure, termed ‘Lamellar’, is demonstrated by Fig. 2(d), where one non-zero cone angle value produces a structure featuring several interconnected layers, with the layer count being proportional to the wave number. Demonstration of controlling the anisotropy of structures through a choice of cone angles can be found in A.1, which also highlights the symmetry in the design space, where the selected values of $\theta$ merely alter the orientation of the topology. Conversely, Fig. 2(c) showcases the formation of a ‘Cubic’ topology through the use of three non-zero cone angles of equivalent magnitude, which produces a structure with identical stiffnesses in all principal directions. Meanwhile, ‘Isotropic’ structures, as shown in Fig. 2(b), emerge from any combination of angles featuring sufficiently large values of cone angles.
@@ 2.2 FEM setup and material characterisation
The spinodoid topologies with dimensions $L\times L\times L$ (where $L=40$ mm) were generated using MATLAB, and FEM simulations were performed using the ABAQUS Explicit solver. Tetrahedral elements were used for meshing, and the structures were compressed along the $x_{3}$ direction. Due to the symmetry of design parameters, $\theta$, compression in a single direction was sufficient to represent the behaviour across the entire design space, as illustrated in A.1. The simulation setup included two rigid plates: one acting as the loader, which moved until a maximum strain of 50% was achieved, and the other serving as a stationary anvil. In addition, general contact was defined to simulate interactions between the platens and the structure as well as self-contact, enabling densification and preventing self-interaction during deformation.
Since the spinodoid topologies were fabricated using additive manufacturing, directional dependencies were inherently introduced. These dependencies arose due to the layer-by-layer printing process, which affects the mechanical response depending on the orientation of the layers relative to the loading direction. Capturing this anisotropy in the simulation required an accurate representation of the material properties. This effect was taken into account by utilising engineering constants to describe the linear elasticity in orthotropic materials. This involved calculating three moduli in each principal direction $E_{1},E_{2},E_{3}$; Poisson’s ratio $\nu_{12},\nu_{13},\nu_{23}$; and the shear moduli $G_{12},G_{13},G_{23}$. These constants are then used to construct the elastic compliance, which is as follows
@@
${\color[rgb]{0,0,0}\sqrt{F(\sigma_{22}-\sigma_{33})^{2}+G(\sigma_{33}-\sigma_{11})^{2}+H(\sigma_{11}-\sigma_{22})^{2}+2L\sigma^{2}_{23}+2M\sigma^{2}_{31}+2N\sigma^{2}_{12}}=f(\mathbf{\sigma})}$ (5)
${\color[rgb]{0,0,0}F=\frac{1}{2}\left[\frac{1}{R_{22}^{2}}+\frac{1}{R_{33}^{2}}-\frac{1}{R_{11}^{2}}\right]\qquad L=\frac{3}{2R^{2}_{23}}}$ (6)
${\color[rgb]{0,0,0}G=\frac{1}{2}\left[\frac{1}{R_{33}^{2}}+\frac{1}{R_{11}^{2}}-\frac{1}{R_{22}^{2}}\right]\qquad M=\frac{3}{2R^{2}_{13}}}$
${\color[rgb]{0,0,0}H=\frac{1}{2}\left[\frac{1}{R_{11}^{2}}+\frac{1}{R_{22}^{2}}-\frac{1}{R_{33}^{2}}\right]\qquad N=\frac{3}{2R^{2}_{12}}}$
where $R_{11}=\frac{\sigma_{11}^{y}}{\sigma^{0}},R_{22}=\frac{\sigma_{22}^{y}}{\sigma^{0}},R_{33}=\frac{\sigma_{33}^{y}}{\sigma^{0}},R_{12}=\frac{\sigma_{12}^{y}}{\tau^{0}},R_{13}=\frac{\sigma_{13}^{y}}{\tau^{0}},R_{23}=\frac{\sigma_{23}^{y}}{\tau^{0}}$, with $\tau_{0}=\frac{\sigma_{0}}{\sqrt{3}}$. $R_{ij}$ represents the anisotropic yield stress ratios, with $\sigma^{y}_{ij}$ being the yield stress values when only $\sigma_{ij}$ has been applied. Here, $\sigma_{0}$ is a user-defined yield stress reference.
@@
To account for these effects, material properties were obtained through material characterisation tests conducted in three directions: along the printed layers, orthogonal to it, and at a $45^{\circ}$ angle in the same plane to obtain the shear behaviour. Tensile and compressive tests were performed on sample type 5A of the ISO-527 standards and $20\times 20\times 20$ $\text{mm}^{3}$ cubes, respectively, with the smaller cube size selected to avoid requiring a high-capacity load cell for compression tests. Both quasi-static compression and tensile tests were conducted at a strain rate of $0.1~{}s^{-1}$ using an INSTRON-5967 machine. Test specimens were fabricated via fused deposition modelling on an UltiMaker S5, printed at three orientations: $0^{\circ}$, $45^{\circ}$, and $90^{\circ}$, representing the directions along the loading axis to obtain material properties. The results of these tests have been graphed in Fig. 3.
Material properties were extracted from the experimental data to define the constitutive model for the simulations. The parameters for the Hill anisotropic plasticity model were determined based on stress-strain data obtained from the material characterisation tests. The Young’s modulus values were assigned as $E_{1}=E_{0^{\circ}}$, reflecting the behaviour in the plane parallel to the build plate, but perpendicular to the loading direction, while $E_{2}=E_{3}=E_{90^{\circ}}$ corresponds to the plane orthogonal to the printing direction, but parallel to the loading direction. Shear moduli were approximated using the relation $G_{12}=G_{13}=G_{23}=E_{45^{\circ}}/2(1+\nu)$, where $E_{45^{\circ}}$ is the modulus obtained from the $45^{\circ}$ sample, with the assumption of uniform Poisson’s ratio across orientations at a value of 0.3. Yield strengths were extracted using the 0.2% offset method: $\sigma_{11}^{y}=\sigma_{0^{\circ}}^{y}$ were obtained for in-plane direction, $\sigma_{22}^{y}=\sigma_{33}^{y}=\sigma_{90^{\circ}}^{y}$ for out of plane direction, and $\sigma_{12}^{y}=\sigma_{45^{\circ}}^{y}$ was inferred from the $45^{\circ}$ orientation for the shear strength to represent shear strength was based on its alignment with the principal material directions in the Hill model. In addition, the reference yield stress to obtain the yield stress ratio was set to be $\sigma_{0}=\sigma_{11}^{y}$. The selection of directions for applying these material properties was based on the orientation of the spinodoid geometry slices used for 3D printing the structure, in this case, the geometries were sliced along the $x_{1}-x_{2}$ plane i.e., printed perpendicular to the direction of loading.
After determining the material parameters required to construct the Hill anisotropic model using the characterisation tests, the accuracy of the fitted parameters was evaluated through an initial validation test. Specifically, FEM simulation was compared with experimental results from one of the material characterisation tests to ensure the validity and robustness of the model. A $90^{\circ}$ cube under compression was selected as the benchmark due to its printing orientation, which closely aligned with the fabricated spinodoid topologies, ensuring a representative comparison. Furthermore, the cube geometry was particularly suitable for this validation because the primary performance assessment of the spinodoid topologies in this study is under compressive loading conditions. The use of this setup allowed a direct and meaningful comparison between simulation and experiment, providing confidence in the accuracy of the material model. While the $90^{\circ}$ cube is not isotropic, its geometry and orientation were specifically chosen to reflect the intended application of the material. This ensures the validation is representative of real-world conditions, particularly the compressive behaviour of spinodoid topologies. The detailed results of this validation test are presented in A.2.
@@ 2.3 Experimental validation
To validate the FEM model, the four spinodoid structures described in Fig. 2 were employed. Similar to the testing samples, the spinodoid structures of dimensions $20\times 20\times 20$ mm3 were sliced using UltiMaker Cura, and ultimately additively manufactured using UltiMaker S5. These prototypes were printed using PET-G, with PVA as the support material, which was later removed by immersing the printed samples in water for an extended duration.
@@
Furthermore, the tests were conducted on samples with a relative density of 0.3 and a wave number of $24\pi$. A smaller dimension of spinodoid structures was chosen, along with low relative density to offset the densification, and by extension, not require a higher capacity load cell. Compressive testing was conducted using a quasi-static strain rate of 0.1 $s^{-1}$. This strain rate was carefully chosen to closely match the conditions of the material characterisation test, minimising the influence of strain-rate variations on the results. To ensure the separation of scales, a larger wave number was utilised as a result of using a smaller-sized structure, wave numbers larger than $24\pi$ would result in geometries with features that cannot be resolved by the 3D printer, even with layer height set to a minimal.
Fig. 4 shows the results of the compressive testing of the spinodoid structures. The compressive tests were conducted on three samples for each type of topology to ensure repeatability as well as to consider the variation in crushing behaviour resulting from manufacturing defects. In the figure, the black lines show the experimental results, while the red lines show the results extracted from the FEM simulations. The material constants used to build the constitutive material model were derived from compressive testing data. This choice was made because the spinodoid prototypes were primarily subjected to compressive loading during their intended applications, ensuring that the mechanical behaviour of PET-G aligns more closely with the working conditions.
However, directly using the material constants obtained from the material characterisation tests caused an overprescription of stiffness within the structures, resulting in a mismatch between experimental and simulation results. This was remedied by reducing the material constants to values that were 20% of the original (i.e. $0.2\cdot E$, where $E$ is the value of the originally calculated material modulus, with a similar penalty being applied to the material’s yield strength). Introducing this artificial reduction in the stiffness constants accounts for any structural defects and artefacts created during the manufacturing of the prototypes, which may weaken the structural integrity of the structures as a result of improper bonding between adjacent layers.
For ‘Lamellar’-type structures, a unique stiffness penalty was applied to account for their distinctive geometry. These structures, defined by layers oriented perpendicular to the loading direction, required a reduction to 60% of the initially calculated material constants (i.e., 0.6$\cdot$ E). This adjustment specifically targeted topologies where $0\leq\theta_{1,2}\leq 5$ and $0\leq\theta_{3}\leq 45$, as they exhibited similar layer-like characteristics. Again, an equivalent penalty was applied to the yield strength for that particular orientation.
The variation in stiffness correction factors across different topologies highlights differences in defect levels caused by the printing orientation. Geometries with their primary features aligned parallel to the print bed generally display fewer defects, as the larger surface area allows subsequent layers to form stronger bonds. In contrast, structures with smaller features are more prone to defects. Additionally, the printing orientation significantly influences the mechanical behaviour of the structures highlighted by the need to use anisotropic material model in FEM simulations. For instance, ‘Columnar’-type structures with columns aligned parallel to the print bed require less reduction in material parameters compared to those with columns oriented perpendicular to the bed. This disparity stems from the printer’s inability to resolve small features, which significantly affects the structural integrity of the spinodoid prototypes unless a nozzle with a very small diameter is utilised. As a result, applying varying stiffness correction factors based on geometry enables the consideration of real-life printing defects.
Nonetheless, the difference between the experiment and the FEM simulations can be a result of a number factors. Firstly, this could be a result of fracture being a pervasive mode of failure, which is not considered in the current material model, where only elasticity and anisotropic plasticity are considered. This is especially apparent in the case of ‘Columnar’-type structure. Additionally, the difference can be caused by the inherent assumptions made when calculating the material parameters for the anisotropic Hill model. This difference can be clearly seen in A.2, whereby the initial elastic strain with elastic constants being directly calculated from the particular stress-strain response can lead to a lower stiffness than initially prescribed, which could be a result of the initial FEM configuration. This could be one of the reasons why there is a small difference in the initial elastic response between the experiment and FEM simulations for the spinodoid structures. Nonetheless, a mixture of mass scaling and time scaling were also utilised to reduce computation time, which can also contribute to the initial non-linear response in the stress-strain response obtained from FEM simulation at small strains, which has been demonstrated in A.6. On the other hand, the disparity in large strain scale (i.e. $\varepsilon>0.4$) for some structures could be a result of reduced porosity in the 3D-printed spinodoid prototypes as a result of manufacturing defects such as stringing, prohibiting water from permeating the structure, allowing the PVA support material to dissolve, resulting in faster densification to occur at high strain values especially in the case of ‘Cubic’ structure due to its unique geometry. Furthermore, the discrepancy between experimental results and FEM simulations can also be attributed to the fact that material parameters were derived solely from compression tests. Given that PET-G exhibits significantly different behaviour under tension and compression, addressing this disparity would require the implementation of advanced material models, such as those defined through User-Defined Material Models (UMAT/VUMAT).